Fermi National Laboratory
 

Accelerator Update for December 21, 2006 - January 3, 2007

 

Thursday December 21

The day shift began with Operations monitoring store 5137 (90.67E30) and a Recycler stash, with the Antiproton Source (Pbar) stacking (25.7E10), with Linac, Booster, Main Injector (MI), and Recycler operating normally, with MiniBooNE and NuMI taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

The day shift began with D0 in access until 10 AM. Technicians worked on a Linac RF station (KRF6) until 11:30 AM. Tevatron (TeV) experts conducted studies until the TeV suffered a sector F1 quench at 3 PM. The Recycler completed one stash of antiprotons.

The TeV studies continued into the evening shift. Operations put the TeV into shot setup and established store 5143 at 11 PM with an initial luminosity of 180.0E30. Recycler experts conducted studies and competed 3 stashes. The I- Source sparked several times.

Friday December 22

The midnight shift began with Operations monitoring store 5143 (136.9E30) and the stash (47.2E10), with Pbar stacking (13.0E10), and with MiniBooNE and NuMI taking beam.

The midnight shift began with a TeV sector B17 separator spark. The Recycler completed 2 stashes.

On the day shift a blown fuse caused a Recycler cooling problem. An expert worked on a Booster RF station (BRF19).

At 11 PM on the evening shift Operations terminated store 5143. D0 accessed their collision hall. The Recycler completed 3 stashes.

Saturday December 23

The midnight shift began with Operations monitoring a stash (325.7E10), with Pbar stacking (46.2E10), with D0 in access, and with NuMI and MiniBooNE taking beam.

D0 came out of access at 1:11 AM on the midnight shift. The TeV quenched during a dry squeeze. No cause was found. Operations successfully put the TeV through a dry squeeze at 4:30 AM. Operations established store 5144 at 7:30 AM with an initial luminosity of 196.8E30. Recycler RF problems held back 80E10 of antiprotons from the store. More RF problems arose during the stash. The Recycler completed 1 stash.

On the day shift, store 5144 was lost due to a sector E2 quench.

The evening shift began with Operations establishing store 5146 at 5:01 PM with an initial luminosity of 170.2E30. The Recycler completed 1 stash.

Sunday December 24

The midnight shift began with Operations monitoring store 5146 and the stash, with Pbar stacking, and with MiniBooNE and NuMI taking beam.

By the end of midnight shift, the Recycler had completed 5 stashes.

On the day shift, Operations terminated store 5146 at 10:12 AM due to a cold box failure. Cryo system experts worked on this through the day and early evening shift. Also during the morning, a Booster kicker failed (MKS07).

Operations established store 5148 at 8:45 PM on the evening shift with an initial luminosity of 221.75E30. The Recycler had completed 3 stashes by the end of the shift.

Monday December 25

The midnight shift began with Operations monitoring store 5148 and a stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

At 2:52 AM on the midnight shift, store 5148 was terminated due to another cold box crash. Operators traced the problem to a faulty UPS. The I- Source sparked several times. The Recycler completed 4 stashes.

On the day shift, Operations established store 5150 at 11 AM with an initial luminosity of 163.37E30. A Linac RF station (KRF3) had cooling system regulation problems. The Recycler completed 2 stashes.

The evening shift ended with the Recycler completing its last of 2 stashes.

Tuesday December 26

The midnight shift began with Operations monitoring store 5150 (42.98E30) and the stash (196.6E10), with Pbar stacking (14.3E10), and with MiniBooNE and NuMI taking beam.

During the midnight shift, MI-10 suffered a vacuum burst. The Recycler completed 3 stashes.

The day shift started with Operations terminating store 5150 at 8:01 AM. A Pbar magnet in the IB4 transport line developed a LCW leak due to a blown hose. Techs repaired the leak. Operations established store 5151 at 10:25 AM with an initial luminosity of 207.0E30. CHL lost one of its big cold compressor. A kicker at F17 tripped off.

Techs completed repairs on the F17 kicker on the evening shift. The Recycler completed 2 stashes.

Wednesday December 27

The midnight shift began with Operations monitoring store 5151 (55.89E30) and a stash (107.3E10), with Pbar stacking (50.2E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, a Linac RF station (KRF2) held off beam due to imbalance faults for about 40 minutes. The Recycler completed 3 stashes.

The Recycler completed one stash during the day shift. Operations terminated store 5151 at 2 PM.

At 5:59 PM on the evening shift, Operations established store 5152 with an initial luminosity of 215.36E30. The Recycler completed 3 stashes.

Thursday December 28

The midnight shift began with Operations monitoring store 5152 (98.35E10) and the stash (127.2E10), with Pbar stacking (19.5E10) and with MiniBooNE and NuMI taking beam.

The I- Source was sparking on the midnight shift. Linac had buncher problems. Operators had to reset the Booster WAPS and two MI anodes. The Recycler stashed twice.

On the day shift a Linac RF cavity (LRF5) pump had a LCW leak, which was fixed. Later, the LRF5 pump burned up and was replaced. Operators switched to the H- Source.

During the evening shift Operators terminated store 5152 and at 11:49 PM established store 5153 with an initial luminosity of 207E30.

Friday December 29

The midnight shift began with Operations monitoring store 5153 (179.61E10) and a stash (54.4E10), with Pbar stacking (20.9E10), and with NuMI and MiniBooNE taking beam.

The midnight shift saw TRF1 trip off; operators reset it. Operators began tuning to resolve TeV lifetime problems.

On the day shift, an RF expert reported that TRF1 was causing the lifetime issues. Operators turned the station off. The Recycler stashed three times.

At 10:02 PM, on the evening shift, operators terminated store 5153. D0 made an access. The Recycler stashed three times.

Saturday December 30

The midnight shift began with Operations monitoring the stash (354.9E10), with Pbar stacking (14.5E10), and with NuMI and MiniBooNE taking beam.

Operations established store 5155 at 3:20 AM on the midnight shift with an initial luminosity of 248.204E30. This store set a NEW RECORD LUMINOSITY. The Recycler stashed twice.

The day shift was quiet except for an air leak on the TeV C4 air compressor. Experts repaired the leak. Operations terminated store 5155 at 9 PM for a D0 access. The Recycler stashed six times.

Sunday December 31

The midnight shift began with the TeV in shot setup, with Operations monitoring the stash, and with MiniBooNE and NuMI taking beam.

On the midnight shift, Operations established store 5157 at 1:49 AM with an initial luminosity of 235.9E30. Operators had to reset LRF5 twice. The Recycler stashed four times.

Pbar problems either kept the stack rate low or halted stacking for most of the day shift. A Pbar Debuncher quadrupole (D:QF) tripped off twice. The Recycler stashed twice.

The evening shift began with Operations turning off MiniBooNE as scheduled for absorber work. Pbar experts investigated their stacking problems, which led to reconfiguring the Debuncher RF. Normal stacking resumed. The Recycler stashed twice.

Monday January 1

The midnight shift and the New Year began with Operations monitoring store 5137 and a stash, with Pbar stacking, and with NuMI taking beam.

The Recycler stashed twice during the midnight shift.

The day shift began with Operations terminating store 5157 at 8:02 AM. Operations established store 5158 at 10:12 AM with an initial luminosity of 243.7E30. Operations turned off NuMI briefly due to a fire alarm at the MINOS far detector. Operators tuned Booster. The Recycler stashed twice.

The Recycler stashed three times during the evening shift.

Tuesday January 2

The midnight shift began with Operations monitoring store 5158 (69.42E30) and the stash (207.6E10), with Pbar stacking (33.7E10), and with NuMI taking beam.

The Recycler stashed three times during the midnight shift.

The day shift began with Operations terminating store 5158 at 8 AM. D0 and CDF accessed their collision hall. Cryo system technicians replaced a sector E3 wet engine flywheel. Operations established store 5160 with an initial luminosity of 225.0E30 at 1:37 PM. A B17 separator spark affected the luminosity. Instrumentation experts worked on installing a new instrument to measure the Recycler beam intensity. The Recycler stashed twice during the shift.

Pbar experts conducted stacktail studies early on the evening shift. The Recycler stashed three times.

Wednesday January 3

The midnight shift began with Operations monitoring store 5160 (78.3E30) and the stash (150.6E10), with Pbar stacking (46.8E10), and with NuMI taking beam.

During the midnight shift, a MI vacuum burst held off beam for ten minutes. The Recycler stashed three times.

The Plans for Wednesday

The plans for today are to stack, stash, and store. The Run Coordinator will allow Linac and Booster experts to conduct studies this afternoon.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 01/03/2007    email Fermilab